OpenStudy (torilucas1):

what are the first four terms of the sequence given by the formula below An=-2n+4 A)-6,-8,-10,-12 B)-2,0,2,4 C)2,0,-2,-4 D)6,8,10,12

OpenStudy (palwandew):

Its option B. Put n=1,2,3,4 in given formula and solve.

OpenStudy (whpalmer4):

@Palwandew might want to check your work there... \[A_n = -2n+4\]As an example, I will find the 5th term:\[A_5 = -2(5)+4 = -10+4 = -6\]

OpenStudy (palwandew):

@whpalmer4 Oops. Sorry its option C.